OpenStudy (anonymous):

45 in fractional notation 45/100 and from there i am supposed to simplify but i always get confused on this part

OpenStudy (anonymous):

What are the common factors that are same from 45 and 100?

jimthompson5910 (jim_thompson5910):

Similar Example Reducing 15/100 The two numbers 15 and 100 have the greatest common factor (GCF) of 5 since this is the largest factor of both 15 and 100. Now you divide each number by the GCF 15/5 = 3 100/5 = 20 That means that 15/100 reduces to 3/20. In other words, 15/100 = 3/20 Note: this is NOT the answer as it's just an example

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome